OpenStudy (anonymous):

Cyrus is shopping for a printer cartridge at the office supply store. Which option is the lowest price? Table showing prices and incentives for different brands - first column Brand, second column Regular price, third column Incentive. First row Brand store brand, Regular price 22 dollars and 95 cents, Incentive none. Second row Brand brand 1, Regular price 33 dollars and 87 cents, Incentive coupon for 20 percent off. Third row Brand brand 2, Regular price 34 dollars and 45 cents, Incentive 10 dollar mail-in rebate. A. store brand B. brand 1 C. brand 2

OpenStudy (texaschic101):

store brand = 22.95 ....no incentive brand 1 = 33.87....20% off brand 2 = 34.45...$ 10 mail in rebate do I have these correctly written according to your chart ??

OpenStudy (anonymous):

YES

OpenStudy (texaschic101):

ok...we already know the store brand (it equals 22.95) brand 1.....33.87 - 20% of 33.87 = 33.87 - 0.20(33.87) = 33.87 - 6.77 = 27.10 brand 2....34.45 - 10 = 33.45 so what is the lowest ??

OpenStudy (anonymous):

BRAND 1

OpenStudy (texaschic101):

check again....

OpenStudy (anonymous):

STORE BRAND

OpenStudy (texaschic101):

yes..thats it....you got it :)

OpenStudy (anonymous):

THX SOO MUCH

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

Nicole is comparing prices for the same brand and model of an electric guitar. Which store offers the lowest price? Store Table showing prices and incentives at different stores - first column Store, second column Regular price, third column Incentive. First row Store A, Regular price 125dollars and 97 cents, Incentive 15 dollar mail-in rebate and 10 dollars off sale. Second row Store B, Regular price 108 dollars and 26 cents, Incentive coupon for 10 percent off a regular-price item. Third row Store C, Regular price 99 dollars and 95 cents, Incentive 5 dollar instant rebate. Fourth row, Store D, Regular price 112 dollars and 8 cents, Incentive sale for 15 percent off.